Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
ciardelm - 14:27 Tuesday 31 March 2020 (48876) Print this report
TCS CO2 laser cooling chillers check (with water refilling to both chillers)
The CO2 laser cooling chillers have been checked this morning (around noon):
- NI chiller has been filled at level
- WI chiller has been filled at level
No other operation was necessary.
Comments to this report:
dattilo - 17:59 Friday 03 April 2020 (48902) Print this report
The CO2 laser cooling chillers have been checked this afternoon (around 15:40 UTC).
No need of refill.
ciardelm - 15:27 Tuesday 07 April 2020 (48909) Print this report
The CO2 laser cooling chillers have been checked this morning (around 10:00 UTC).
No need of refill.
berni - 16:51 Friday 10 April 2020 (48916) Print this report

The CO2 laser cooling chillers have been checked and refilled today around 13:00 UTC.

ciardelm - 19:17 Tuesday 14 April 2020 (48919) Print this report
The CO2 laser cooling chillers have been checked this morning (07h06 UTC):
- NI chiller has been filled at level
- WI chiller has been filled at level
No other operation was necessary.
ciardelm - 15:56 Tuesday 21 April 2020 (48922) Print this report
The CO2 laser cooling chillers have been checked at 12:54 UTC.
No need of refill.
ciardelm - 16:06 Friday 24 April 2020 (48931) Print this report
The CO2 laser cooling chillers have been checked this afternoon (13h00 UTC):
- WI chiller has been filled at level
No other operation was necessary.
ciardelm - 17:37 Tuesday 28 April 2020 (48935) Print this report
The CO2 laser cooling chillers have been checked this afternoon (12h54 UTC):
- NI chiller has been filled at level
No other operation was necessary.
ciardelm - 15:33 Tuesday 05 May 2020 (48942) Print this report
The CO2 laser cooling chillers have been checked this morning (07h54 UTC):
- NI chiller filled at level
- WI chiller filled at level
In addition, the bench utilities cooling system water filter was replaced.
(QC check: no cooling water leakage)
ciardelm - 10:18 Tuesday 12 May 2020 (48944) Print this report
The CO2 laser cooling chillers have been checked this morning (06h40 UTC):
- NI chiller filled at level
- WI chiller filled at level
No other operation was necessary.
berni - 10:10 Friday 15 May 2020 (48946) Print this report

The chillers have been checked this morning; everything was working properly.

ciardelm - 15:14 Tuesday 19 May 2020 (48952) Print this report
The CO2 laser cooling chillers have been checked this morning (10h07 UTC):
- NI chiller filled at level
- WI chiller filled at level
No other operation was necessary.
ciardelm - 12:02 Tuesday 26 May 2020 (48960) Print this report
The CO2 laser cooling chillers have been checked this morning (09h45 UTC):
- WI chiller filled at level
No other operation was necessary.
Search Help
×

Warning

×